). Use the coordinates (1997, 1.26) and (2006, 2.31) to find the slope, or rate of change, between the two points. Describe how you arrived at your answer.

Answer by lwsshak3(11628)

Use the coordinates (1997, 1.26) and (2006, 2.31) to find the slope, or rate of change, between the two points. Describe how you arrived at your answer.
**
Slope is defined as the change in y (dependent variable)/change in x (independent variable)
For given ordered pairs, the first number is x, and the second y.
slope=∆y/∆x=(2.31-1.26)/(2006-1997)=1.05/9